OpenStudy (anonymous):

suppose that cos(t)=2/3. Find the exact value of sec(t). Help please.

OpenStudy (anonymous):

sec(t) = 1/(cos)t = 1 / 2/3 = 3/2

OpenStudy (jamesj):

Right. By definition\[\sec t = \frac{1}{\cos t}\]

OpenStudy (anonymous):

Wow, this is awesome. Thank you both so much.. :)

OpenStudy (anonymous):

yw